Find the asymptotes - x3 - 2x2y - 2x2 + 4xy - 4y2 + 4x + y + 3 = 0
1
Expert's answer
2020-12-29T17:16:43-0500

The given equation of the curve is "-x^3-2x^2y-2x^2+4xy-4y^2+4x+y+3=0"

We shall first find the vertical asymptotes of the if any :

The highest power of the term of "y" is "y^2." Its coefficient is "-4" (constant).

So the curve has no vertical asymptotes.

Next we shall find horizontal asymptote:

Let "y=mx+c" be the horizontal asymptote.

Then "m" and "c" can be find in the following way.

We arrange the equation of the curve as

"(-x^3-2x^2y)+(-2x^2+4xy-4y^2)+(4x+y)+3=0"

"\implies x^3[-1-2.\frac{y}{x}]+x^2[-2+4.\frac{y}{x}-4(\frac{y}{x})^2]+(4x+y)+3=0"


Here , "\phi_3(\frac{y}{x})=-1-2(\frac{y}{x})"

"\phi_2(\frac{y}{x})=-2+4(\frac{y}{x})-4(\frac{y}{x})^2"

"\phi _1(\frac{y}{x})=4x+y"

"\phi_0(\frac{y}{x})=3"

Now consider the equation, "\phi_3(m)=0"

"\implies-1-2m=0"

"\implies m=-\frac {1}{2}"

Again we know that , "c=-\frac{\phi_2(m)}{\phi_3'(m)}" [ if "\phi_3'(m)\neq0" ]

Here "\phi_3'(m)=-2" and "\phi_2(m)=-2+4m-4m^2"

Now corresponding to "m=-\frac{1}{2}" , "c=-\frac{\phi_2(-\frac{1}{2})}{\phi_3'(-\frac{1}{2})}"

"\therefore \phi_2(-\frac{1}{2})=-5" and "\phi_3'(-\frac{1}{2})=-2" and "c=-\frac{5}{2}"

Hence "y=(-\frac{1}{2}).x+(-\frac{5}{2})" is an asymptote.

Therefore "x+2y+5=0" is the only horizontal asymptote of the given curve.